Suppose the weather of the day for the future days (starting from tomorrow) is good (including sunny, cloudy and etc) has probability 1/3 and the weather of the day is bad (including raining, foggy, windy and etc.) has probability 2/3.

Starting from tomorrow, compute the expectation of number of days until there are 7 consecutive days of good weather.
Answer #1

We are given here that:
P( good weather) = 1/3 and P( bad weather ) = 2/3

Let the expected number of days it takes to reach 7 consecutive days of good weather be X.

Let the expected number of days be A1 when 1 good weather day has already occurred.
Let the expected number of days be A2 when 2 good weather day has already occurred.
Let the expected number of days be A3 when 3 good weather day has already occurred.
Let the expected number of days be A4 when 4 good weather day has already occurred.
Let the expected number of days be A5 when 5 good weather day has already occurred.
Let the expected number of days be A6 when 6 good weather day has already occurred.

Therefore, X = (1/3)(A1 + 1) + (2/3)*(X + 1)
3X = A1 + 1 + 2X + 2
X = A1 + 3

From 1 good weather already occurred, we have here:
A1 = (2/3)*(X + 1) + (1/3)*(A2 + 1)
3A1 = 2X + 2 + A2 + 1
3A1 = 2A1 + 6 + 2 + A2 + 1
A1 = A2 + 9

From 2 good weather already occurred, we have here:
A2 = (2/3)*(X + 1) + (1/3)*(A3 + 1)
3A2 = 2X + 2 + A3 + 1
3A1 - 27 = 2A1 + 6 + 2 + A3 + 1
A1 = A3 + 36

From 3 good weather already occurred, we have here:
A3 = (2/3)*(X + 1) + (1/3)*(A4 + 1)
3A3 = 2X + 2 + A4 + 1
3(A1 - 36) = 2A1 + 6 + 2 + A4 + 1
A1 = 108 + 9 + A4
A1 = A4 + 117

and so on.. every time we are multiplying the number by 3 and adding 9 here.
A1 = A5 + 3*117 + 9
A1 = A5 + 360

A1 = A6 + 3*360 + 9
A1 = A6 + 1089

A6 = (2/3)(X + 1) + (1/3)*1
3A6 = 2X + 2 + 1
3A1 - 3267 = 2A1 + 6 + 3
A1 = 3276

Therefore, X = 3276 + 3 = 3279

Therefore 3279 is the expected number of days until there are 7 consecutive days of good weather.